OpenStudy (anonymous):

The product of a number n and 2n-5 is equal to 250. What should be the value of n?

OpenStudy (anonymous):

n(2n-5)=250 2n^2-5n-250=0 n=25/2 or n=-10

OpenStudy (anonymous):

we can check it by putting back n=25/2 or n=-10 in it